# Mean-field underdamped Langevin dynamics and its space-time discretization

### 01 Jan 2024 v2

@@ Appendix C Continuous-time results
$\leq(\mathcal{E}^{N}(\mu_{0}^{N})-N\mathcal{E}(\mu_{*}))\exp\left(-\frac{\mathscr{C}_{\textsf{LSI}}}{24\mathscr{L}}t\right)+\frac{\mathscr{L}dt}{\mathscr{C}_{\textsf{LSI}}}(5\mathscr{C}_{\textsf{LSI}}+3\mathscr{L})\exp\left(-\frac{\mathscr{C}_{\textsf{LSI}}}{24\mathscr{L}}t\right)$
@@ Appendix D Discretization analysis
$=\mathbb{E}_{\mathbf{Q}_{Kh}}\sum_{k=0}^{K-1}\left(-\frac{1}{\sqrt{2\gamma}}\int_{kh}^{(k+1)h}\left\langle\left(\begin{array}[]{c}0\\ {D}_{\mu}F(\mu_{t}^{x},x_{t})-{D}_{\mu}F(\mu_{kh}^{x},x_{kh})\end{array}\right),\mathrm{d}\mathrm{B}_{t}\right\rangle\right.$
@@
$\leq\frac{{\mathscr{L}}^{2}}{2\gamma}\sum_{k=0}^{K-1}\int_{kh}^{(k+1)h}\mathbb{E}_{\mathbf{Q}_{Kh}}\|x_{t}-x_{kh}\|^{2}+W_{2}^{2}(\mu_{t}^{x},\mu_{kh}^{x})\mathrm{d}t$
@@
$\lesssim\sqrt{\textsf{KL}(\mathbf{Q}_{Nh}\|\mathbf{p}_{Nh})}+\sqrt{\textsf{KL}(\mu_{Kh}\|\mu_{*})}$
@@
$\leq\frac{\mathscr{L}^{2}}{2\gamma}\sum_{k=0}^{K-1}\int_{kh}^{(k+1)h}\frac{1}{N}\sum_{i=1}^{N}\mathbb{E}_{\mathbf{Q}^{i}_{Kh}}\|x^{i}_{t}-x^{i}_{kh}\|^{2}+W_{2}^{2}({\mu_{\textbf{x}_{t}}},\mu_{\textbf{x}_{kh}})\mathrm{d}t$
@@ Appendix F Methods for comparisons
where $\delta_{hk}\triangleq\mathbb{E}\|{D}_{\mu}F(\bar{\mu}_{k+1},x_{k+1})-{D}_{\mu}F(\bar{\mu}_{k},x_{k})\|^{2}$ and $\mathbb{E}$ is taken under the joint law of $\bar{\mu}_{k+1}$ and $\bar{\mu}_{k}$. Now we bound $\delta_{hk}$ uniformly in $k$ with a different method from the one in Nitanda et al. (2022). We do not need to specify $F$ to be the objective of training nerual networks. Since $F$ is $\mathscr{L}$-smooth and satisfies Assumption 2.4, we obtain

### 17 Jan 2024 v3

@@ Appendix C Continuous-time results
$\leq(\mathcal{E}^{N}(\mu_{0}^{N})-N\mathcal{E}(\mu_{*}))\exp\left(-\frac{\mathscr{C}_{\textsf{LSI}}}{24\mathscr{L}}t\right)+\frac{\mathscr{L}\mathrm{d}t}{\mathscr{C}_{\textsf{LSI}}}(5\mathscr{C}_{\textsf{LSI}}+3\mathscr{L})\exp\left(-\frac{\mathscr{C}_{\textsf{LSI}}}{24\mathscr{L}}t\right)$
@@ Appendix D Discretization analysis
$=\mathbb{E}_{\mathbf{Q}_{Kh}}\sum_{k=0}^{K-1}\left(-\frac{1}{\sqrt{2\gamma}}\int_{kh}^{(k+1)h}\left\langle\left(\begin{array}[]{c}0\\ {D}_{\mu}F(\mu_{t}^{x},x_{t})-{D}_{\mu}F(\mu_{kh}^{x},x_{kh})\end{array}\right),\mathrm{dB}_{t}\right\rangle\right.$
@@
$\leq\frac{{\mathscr{L}}^{2}}{2\gamma}\sum_{k=0}^{K-1}\int_{kh}^{(k+1)h}\mathbb{E}_{\mathbf{Q}_{Kh}}\|x_{t}-x_{kh}\|^{2}+W_{1}^{2}(\mu_{t}^{x},\mu_{kh}^{x})\mathrm{d}t$
@@
$\lesssim\sqrt{\textsf{KL}(\mathbf{Q}_{Kh}\|\mathbf{p}_{Kh})}+\sqrt{\textsf{KL}(\mu_{Kh}\|\mu_{*})}$
@@
$\leq\frac{\mathscr{L}^{2}}{2\gamma}\sum_{k=0}^{K-1}\int_{kh}^{(k+1)h}\frac{1}{N}\sum_{i=1}^{N}\mathbb{E}_{\mathbf{Q}^{i}_{Kh}}\|x^{i}_{t}-x^{i}_{kh}\|^{2}+W_{1}^{2}({\mu_{\textbf{x}_{t}}},\mu_{\textbf{x}_{kh}})\mathrm{d}t$
@@ Appendix F Methods for comparisons
where $\delta_{hk}\triangleq\mathbb{E}\|{D}_{\mu}F(\bar{\mu}_{k+1},x_{k+1})-{D}_{\mu}F(\bar{\mu}_{k},x_{k})\|^{2}$ and $\mathbb{E}$ is taken under the joint law of $\bar{\mu}_{k+1}$ and $\bar{\mu}_{k}$. Now we bound $\delta_{hk}$ uniformly in $k$ with a different method from the one in Nitanda et al. (2022). We do not need to specify $F$ to be the objective of training nerual networks. Since $F$ is $\mathscr{L}$-smooth [^3] We inherit the weaker smoothness assumption in Suzuki et al. (2023) with respect to $W_{2}$ distance. and satisfies Assumption 2.4, we obtain